OpenStudy (anonymous):

36+(-17)+10+(-27) add

OpenStudy (anonymous):

This is the same as 36-17+10-27. :)

OpenStudy (anonymous):

36 -17+10-27 = 9 + 10 -27 = 19 -27 = -8

OpenStudy (anonymous):

Bingo. :)

OpenStudy (anonymous):

what is my answer